# GLHF: General Learned Evolutionary Algorithm Via Hyper Functions

### v1

@@ Abstract
Pretrained Optimization Models (POMs) leverage knowledge gained from optimizing various tasks, providing efficient solutions for new optimization challenges through direct usage or fine-tuning. Despite the inefficiencies and limited generalization abilities observed in current POMs, our proposed model, the general pre-trained optimization model (GPOM), addresses these shortcomings. GPOM constructs a population-based pretrained Black-Box Optimization (BBO) model tailored for continuous optimization. Evaluation on the BBOB benchmark and two robot control tasks demonstrates that GPOM outperforms other pretrained BBO models significantly, especially for high-dimensional tasks. Its direct optimization performance exceeds that of state-of-the-art evolutionary algorithms and POMs. Furthermore, GPOM exhibits robust generalization capabilities across diverse task distributions, dimensions, population sizes, and optimization horizons.
Pretrained Optimization Models, Evolutionary Strategy, Black-box Optimization.
@@ 1 Introduction
A Black Box Optimization (BBO) algorithm is designed to address black box optimization problems, including tasks like hyperparameter optimization (HPO) (Hutter et al., 2019), neuroevolution (Such et al., 2017), neural architecture search (NAS) (Ye et al., 2021), multi-objective optimization (Deb et al., 2002), and multi-task optimization (Bali et al., 2019). In these scenarios, the algorithm can evaluate $f(\mathbf{x})$ for any solution $\mathbf{x}$; however, access to additional information about $f$, such as the Hessian and gradients, is unavailable.
In light of the "No Free Lunch Theorem" (Wolpert & Macready, 1997), it is recognized that no single algorithm can universally excel across all scenarios. Therefore, addressing diverse BBO problems necessitates the tailored design of specific algorithms to achieve satisfactory performance. Crafting these algorithms typically demands substantial expertise. Presently, researchers heavily rely on manual engineering for BBO algorithm design (Liu et al., 2023b). In this manual design paradigm, the process involves "trial and error" guided by past "optimization experience" (Gämperle et al., 2002; Ronkkonen et al., 2005; Tvrdık, 2006). However, this approach is cumbersome, and the subjective nature of "optimization experience" introduces variability across individuals, posing a risk of domain overfitting. Furthermore, the resulting optimization strategy tends to exhibit limited generalization ability when applied to new tasks.
In recent years, pretrained optimization models, a pivotal facet of AutoML (Hutter et al., 2019), have shown promise in overcoming the limitations of human-designed BBO algorithms. These models, general-purpose optimizers learning from the experience of optimizing multiple tasks, can be applied directly or fine-tuned to efficiently address new optimization tasks. (Chen et al., 2022; Krishnamoorthy et al., 2022, 2023) employ Transformer or diffusion models to pretrain model-based optimizers using offline datasets. While effective, these methods primarily fit optimization trajectories of other BBO algorithms to a specific task, potentially requiring retraining for new tasks, limiting their generalizability. Subsequently, (Lange et al., 2023b) introduced two bilevel optimization frameworks for meta-learning evolution strategy (ES) and genetic algorithm (GA). Furthermore, they still grapple with the curse of dimensionality (Chen et al., 2015) attributed to nonconvex objectives in the outer loop and the constrained search strategy derived from optimization trajectories.
@@
To address the inefficiencies and low generality of existing methods, we introduce the General Pretrained Optimization Model, called GPOM. GPOM constructs a population-based pretrained BBO model for continuous optimization. Leveraging multiple individuals, population-based optimizers gain a better understanding of the fitness landscape. Consequently, employing a population-based Differential Evolution (DE) (Storn & Price, 1997; Thangaraj et al., 2009) mechanism with crossover, mutation, and selection operators, we design Learnable Mutation Module (LMM), Learnable Crossover Module (LCM), and Selection Module for GPOM. These modules equip GPOM with advanced optimization capabilities. Pretraining GPOM on a set of training functions ensures good convergence and generalization. Drawing inspiration from (Finn et al., 2017), we introduce an end-to-end gradient-based training method for GPOM, termed MetaGBT (Meta Gradient-Based Training), ensuring stable and rapid training for GPOM.
We evaluate the performance and and generalization capabilities of GPOM on BBOB (Finck et al., 2010) and two robot control tasks (Brockman et al., 2016). Notably, GPOM exhibits superior optimization capabilities compared to other pretrained BBO models, outperforming them significantly. Even in the context of high-dimensional tasks (4149 dimensions), GPOM excels in directly optimizing new tasks, surpassing state-of-the-art population-based optimizers. Furthermore, GPOM showcases adaptive revision of its optimization strategy (LMM and LCM) based on the current population status, striking a commendable balance between exploration and exploitation. An in-depth analysis reveals that training GPOM on simpler tasks yields better results. Our contributions can be summarized as follows:
  • GPOM stands out as the current leading pretrained BBO model, demonstrating a substantial performance advantage over other models. Its effectiveness extends to direct optimization of new tasks, surpassing state-of-the-art population-based optimizers, even in high-dimensional scenarios.
  • GPOM exhibits remarkable generalization across diverse task distributions, dimensions, population sizes, and optimization horizons, addressing challenges faced by existing POMs.
@@ 2 Related Work
Numerous metaheuristic population-based algorithms, such as genetic algorithms (Holland, 1992), evolution strategies (Hansen & Ostermeier, 2001; Hansen et al., 2003; Ros & Hansen, 2008), particle swarm optimization (Kennedy & Eberhart, 1995; Gong et al., 2015), and differential evolution (Storn & Price, 1997; Stanovov et al., 2022), have been devised to address optimization problems. Notably, CMA-ES (Hansen, 2016) and L-SHADE (Tanabe & Fukunaga, 2014) stand out as state-of-the-art methods for BBO. However, these approaches rely on manually designed components, exhibiting inefficiency and fragility when confronted with new tasks. In contrast, the proposed GPOM can autonomously acquire optimization strategies from problem instances, mitigating the aforementioned limitations.
Pretrained Population-based BBO Algorithms
@@
Pre-training BBO algorithms can be categorized into two types within the meta-learning framework. The first type frames meta-learning BBO algorithms as a bi-level optimization problem (Liu et al., 2022). For instance, (Gomes et al., 2021) leverages meta-learning to infer population-based black-box optimizers that automatically adapt to specific task classes. LES (Lange et al., 2023b) designs a self-attention-based search strategy for discovering effective update rules for evolution strategies through meta-learning. Subsequent works like LGA (Lange et al., 2023a) utilize this framework to discover the update rules of Gaussian genetic algorithms via Open-ES (Zhang et al., 2017). The second type models the meta-learning of a BBO algorithm as a reinforcement learning problem. (Shala et al., 2020) meta-learn a policy that adjusts the mutation step-size parameters of CMA-ES (Hansen, 2016). Category one faces the curse of dimensionality, where an escalating number of optimization strategy parameters leads to skyrocketing training difficulty, impeding the development of intricate strategies. In contrast, category two, which models meta-learning optimizers as reinforcement learning tasks, grapples with training instability. GPOM, employing a gradient-based end-to-end training approach, successfully bypasses the curse of dimensionality, ensuring stable training.
LLM for Optimization
@@
In line with POMs, various optimization approaches leveraging Large Language Models (LLMs) have emerged to address diverse problem domains, including NP-hard problems (Romera-Paredes et al., 2023; Meyerson et al., 2023), algorithm evolution (Liu et al., 2023a; Yang et al., 2023; Lehman et al., 2023), reward design (Ma et al., 2023), and Neural Architecture Search (NAS) (Chen et al., 2023; Nasir et al., 2023). Notably, LLMs play a role in sampling new solutions. However, their optimization strategies depend on externally introduced natural selection mechanisms and are less effective in numerical optimization scenarios. These methods lack the universal applicability as pretrained BBO models due to a deficiency in generating capabilities across tasks.
3 GPOM
@@ 3.1 Problem Definition
A black-box optimization problem can be transformed as a minimization problem, as shown in Eq. (1), and constraints may exist for corresponding solutions:
$\min_{\mathbf{x}}\ f(\mathbf{x}),s.t.\ x_{i}\in[l_{i},u_{i}]$ (1)
where $\mathbf{x}=(x_{1},x_{2},\cdots,x_{d})$ represents the solution of optimization problem $f$, the lower and upper bounds $\mathbf{l}=(l_{1},l_{2},\cdots,l_{d})$ and $\mathbf{u}=(u_{1},u_{2},\cdots,u_{d})$, and $d$ is the dimension of $\mathbf{x}$. A population consists of $n$ individuals, denoted as $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{n}\}$. In this paper, $\mathbf{X}$ is also treated as $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{n}]^{T}$ to support matrix operations. We feed GPOM an initial random population $\mathbf{X}^{0}$ at step 0, specify the evolution generation $T$ for it, and hope that it can generate a population $\mathbf{X}^{T}$ close to the global optimum at step $T$, as shown in Eq. (2).
$\mathbf{X}^{T}=GPOM(\mathbf{X}^{0},T|\theta)$ (2)
$\theta\in\Omega$ is the parameters of GPOM, where $\Omega$ stands for the strategy space. The goal of training GPOM is to find an optimal $\theta$ in $\Omega$.
3.2 Parameterization of DE
DE (Storn & Price, 1997; Thangaraj et al., 2009) constitutes a pivotal algorithmic family within BBO, celebrated for its commendable attributes like rapid convergence and robustness (Das et al., 2016; Neri & Tirronen, 2010).
Mutation Strategy Generalization
Illustratively, considering the classic DE mutation strategy, DE/rand/1, it is defined as $\mathbf{v}_{i}^{t}=\mathbf{x}_{r1}^{t}+F\cdot(\mathbf{x}_{r2}^{t}-\mathbf{x}_{r3}^{t})$ (refer to Appendix A for additional mutation strategies). Here, the integer index $r1$ (likewise, $r2$ and $r3$) is randomly chosen from the range $[0,N]$ and $F$ is constant. The generalized formulation of the mutation strategy is presented as:
Further, we organize the mutation strategy into a matrix form, as shown in Eq. (4).
@@ 3.2 Parameterization of DE
$\mathbf{X}^{t}=[\mathbf{x}_{1}^{t},\mathbf{x}_{2}^{t},\cdots,\mathbf{x}_{N}^{t}],(\mathbf{X}^{t}\in\mathbb{R}^{N\times d})$ is the population in generation $t$ and $\mathbf{S}^{t}$ is the matrix shown as follows:
$\mathbf{S}^{t}=\begin{bmatrix}s_{1,1}^{t}&\cdots&s_{1,N}^{t}\\ \vdots&\ddots&\vdots\\ s_{N,1}^{t}&\cdots&s_{N,N}^{t}\\ \end{bmatrix}$ (5)
$\mathbf{S}$ evolves with each change in $t$, signifying a mutation strategy that adapts across generations. Consequently, it is imperative to devise a module that leverages information from the population at generation $t$ to generate $\mathbf{S}^{t}$.
Crossover Strategy Generalization
3.3 Design of GPOM
We design LMM and LCM to achieve the generation of mutation strategy (that is, generate $\mathbf{S}^{t}$) and crossover strategy (that is, generate $\mathbf{cr}^{t}$), respectively.
LMM
The function of LMM is designed based on Multi-head self-attention (MSA) (Vaswani et al., 2017), as shown as follows:
@@ 3.3 Design of GPOM
$\hat{\mathbf{H}}^{t}$ $=Tanh(\mathbf{H}^{t}\times\mathbf{W}_{m1}+\mathbf{b}_{m1})$ (10)
$\mathbf{Q}^{t}$ $=Tanh(\hat{\mathbf{H}}^{t}\times\mathbf{W}_{m2}+\mathbf{b}_{m2})$
$\mathbf{K}^{t}$ $=Tanh(\hat{\mathbf{H}}^{t}\times\mathbf{W}_{m3}+\mathbf{b}_{m3})$
$\hat{\mathbf{S}}^{t}$ $=Tanh(\frac{\mathbf{Q}^{t}\times(\mathbf{K}^{t})^{T}}{sqrt(d_{m})})$
@@
The topological structure of the population significantly influences their information exchange (Yu et al., 2021). When all individuals engage in information exchange, the algorithm’s convergence may suffer, diversity could diminish, and susceptibility to local optima increases. To address this, we introduce a mask operation during both training and testing phases, where the probability of setting each element in $\hat{\mathbf{S}}^{t}$ to 0 is $r_{mask}$. This operation enhances GPOM’s ability to learn efficient and robust strategies, as validated in our experiments. Consequently, $\mathbf{S}^{t}$ is derived using Eq. (11).
@@
LCM is designed based on FFN (Vaswani et al., 2017), as shown in Eq. (12),
@@
where $\theta_{2}=\{\mathbf{W}_{c1},\mathbf{b}_{c1},\mathbf{W}_{c2},\mathbf{b}_{c2},\mathbf{\tau}\}$ is the parameter of LCM and $\mathbf{Z}^{t}\in\mathbb{R}^{N\times 4}$ is the population information used by LCM. Here, $\mathbf{Z}^{t}=[\mathbf{z}_{1}^{t},\mathbf{z}_{2}^{t},\cdots,\mathbf{z}_{N}^{t}]$. $\mathbf{z}_{i}^{t}$ represents the relevant information of individual $\mathbf{x}^{t}_{i}$ and $\mathbf{X}^{t}$. For example, it can include the ranking information of $\mathbf{x}^{t}_{i}$, the fitness information of $\mathbf{x}^{t}_{i}$, the Euclidean distance between $\mathbf{x}^{t}_{i}$ and $\mathbf{V}^{t}_{i}$, and the distribution information of individuals within the population (such as the fitness distribution, the distance between pairs of individuals), etc. In this paper, $\mathbf{z}_{i}^{t}$ includes the following information as a case study: 1) $\hat{f}_{i}^{t}$: the normalized fitness $f(\mathbf{x}_{i}^{t})$ of $\mathbf{x}_{i}^{t}$; 2) $\hat{r}_{i}^{t}$: the centralized ranking of $\mathbf{x}_{i}^{t}$; 3) $\widehat{fv}_{i}^{t}$: the normalized fitness $f(\mathbf{v}_{i}^{t})$ of $\mathbf{v}_{i}^{t}$; 4) $\widehat{rv}_{i}^{t}$: the centralized ranking of $\mathbf{v}_{i}^{t}$. Among them, $\widehat{fv}_{i}^{t}$ and $\widehat{rv}_{i}^{t}$ are calculated in the same way as $\hat{f}_{i}^{t}$ and $\hat{r}_{i}^{t}$ respectively. The detailed mechanism of LCM is shown in Eq. (13).
$\mathbf{h}^{t}$ $=Tanh(\mathbf{Z}^{t}\times\mathbf{W_{c1}}+\mathbf{b_{c1}})$ (13)
$\mathbf{\hat{h}^{t}}$ $=layernorm(\mathbf{h}^{t}|\mathbf{\tau})$
$\mathbf{cr}^{t}$ $=Sigmoid(\mathbf{\hat{h}^{t}}\times\mathbf{W_{c2}}+\mathbf{b_{c2}})$
where the activation function Sigmoid maps inputs to the range $(0,1)$. $\mathbf{W_{c1}}\in\mathbb{R}^{4\times d_{c}}$, $\mathbf{W_{c2}}\in\mathbb{R}^{d_{c}\times 1}$, $\mathbf{\tau}$ is the learnable parameters of layernorm (Xu et al., 2019). $\mathbf{b_{c1}}$ and $\mathbf{b_{c2}}$ are vectors with dimensions $d_{c}$ and $1$, respectively.
Although we derive $\mathbf{cr}^{t}$ from Eq. (13) as in Eq. (6), the discrete nature of the crossover operator renders it non-differentiable, impeding gradient-based training of the LCM module. To address this limitation, we introduce the gumbel_softmax method (Jang et al., 2016), providing an efficient gradient estimator that replaces non-differentiable samples from a categorical distribution with differentiable samples from a novel Gumbel-Softmax distribution.
@@
$\mathbf{r}_{i}^{t}$ $=rand(d)$ (14)
$\mathbf{cv}_{i}^{t}$ $=gumbel\_softmax(cat(\mathbf{r}_{i}^{t},tile(cr_{i}^{t},d)))$
$\mathbf{u}_{i}^{t}$ $=\mathbf{cv}_{i,0}^{t}\cdot\mathbf{x}_{i}^{t}+\mathbf{cv}_{i,1}^{t}\cdot\mathbf{v}_{i}^{t}$
$\mathbf{U}^{t}$ $=[\mathbf{u}_{1}^{t},\mathbf{u}_{2}^{t},\cdots,\mathbf{u}_{N}^{t}]$
@@
Figure 1: The overall architecture of GPOM.

### v2

# Pretrained Optimization Model for Zero-Shot Black Box Optimization

@@ Abstract
Zero-shot optimization involves optimizing a target task that was not seen during training, aiming to provide the optimal solution without or with minimal adjustments to the optimizer. It is crucial to ensure reliable and robust performance in various applications. Current optimizers often struggle with zero-shot optimization and require intricate hyperparameter tuning to adapt to new tasks. To address this, we propose a Pretrained Optimization Model (POM) that leverages knowledge gained from optimizing diverse tasks, offering efficient solutions to zero-shot optimization through direct application or fine-tuning with few-shot samples. Evaluation on the BBOB benchmark and two robot control tasks demonstrates that POM outperforms state-of-the-art black-box optimization methods, especially for high-dimensional tasks. Fine-tuning POM with a small number of samples and budget yields significant performance improvements. Moreover, POM demonstrates robust generalization across diverse task distributions, dimensions, population sizes, and optimization horizons. For code implementation, see https://github.com/ninja-wm/POM/.
@@ 1 Introduction
Black box optimization, including tasks like hyperparameter optimization (HPO) [1], neuroevolution [2, 3, 4], neural architecture search (NAS) [5], and algorithm selection [6], is very important. In these scenarios, the algorithm can evaluate $f(\mathbf{x})$ for any solution $\mathbf{x}$; however, access to additional information about $f$, such as the Hessian and gradients, is unavailable.
Addressing diverse BBO problems necessitates the tailored design of specific algorithms to achieve satisfactory performance. Crafting these algorithms typically demands substantial expertise. Therefore, it is crucial to ensure reliable and robust performance of the optimizer in various applications, called zero-shot optimization. Zero-shot optimization involves optimizing a target task that was not seen during training, aiming to provide the optimal solution without or with minimal adjustments to the optimizer.
The studies [7, 8, 9] employed Transformer or diffusion models to pretrain model-based optimizers using offline datasets. While effective, these methods primarily fit optimization trajectories of other BBO algorithms to a specific task, potentially requiring retraining for new tasks, limiting their ability to zero-shot optimization. Subsequently, [10, 11] introduced two learned optimization frameworks for meta-learning evolution strategy (ES) and genetic algorithm (GA). However, the performance of these two methods on zero-shot optimization is weaker than that of CMA-ES [12] (see Section 4.2).
@@
To address zero-shot optimization, especially for continuous optimization, we introduce a population-based Pretrained Optimization Model, called POM. Leveraging multiple individuals, population-based optimizers gain a better understanding of the fitness landscape. The core of the optimizer is how to design optimization strategies that sample better solutions. Inspired by the solution-producing mechanism of evolutionary computation, we design powerful POM blocks to form a general optimization strategy representation framework. Drawing inspiration from [13], we introduce an end-to-end gradient-based training method for POM, termed MetaGBT (Meta Gradient-Based Training), ensuring stable and rapid training for POM. Pretraining POM on a set of training functions with MetaGBT ensures good optimization strategy. Our contributions can be summarized as follows:
  • Excellent ability to solve zero-shot BBO. We develop a efficient POM for zero-shot BBO, demonstrating a substantial performance advantage over state-of-the-art black-box optimizers.
  • Excellent ability to solve few-shot BBO. Few-shot optimization is the existence of a small budget of function evaluations for the target task to tune the optimizer for better performance. More than 30% performance improvement can be obtained with 25 random function evaluations.
@@ 2 Related Work
Heuristic Population-based BBO Algorithms. Numerous metaheuristic population-based algorithms, such as genetic algorithms [14], evolution strategies [15, 16, 17], particle swarm optimization [18, 19], and differential evolution [20, 21], have been devised to address optimization problems. Notably, CMA-ES [12] and L-SHADE [22] stand out as state-of-the-art methods for BBO. However, these approaches rely on manually designed components, exhibiting inefficiency and fragility when confronted with new tasks. In contrast, the proposed POM can autonomously acquire optimization strategies from problem instances, mitigating the aforementioned limitations.
@@
Pretrained Population-based BBO Algorithms. Pre-training BBO algorithms can be categorized into two types within the meta-learning framework. The first type frames meta-learning BBO algorithms as a bi-level optimization problem [23]. For instance, [24] leverages meta-learning to infer population-based black-box optimizers that automatically adapt to specific task classes. LES [25] designs a self-attention-based search strategy for discovering effective update rules for evolution strategies through meta-learning. Subsequent works like LGA [10] utilize this framework to discover the update rules of Gaussian genetic algorithms via Open-ES [26]. The second type models the meta-learning of a BBO algorithm as a reinforcement learning problem. [27] meta-learn a policy that adjusts the mutation step-size parameters of CMA-ES [12]. Category one faces the curse of dimensionality, where an escalating number of model parameters leads to skyrocketing training difficulty, impeding the development of intricate strategies. In contrast, category two, which models meta-learning optimizers as reinforcement learning tasks, grapples with training instability. POM, employing a gradient-based end-to-end training approach, successfully bypasses the curse of dimensionality, ensuring stable training.
@@
LLM for Optimization. In line with POMs, various optimization approaches leveraging Large Language Models (LLMs) have emerged to address diverse problem domains, including NP-hard problems [28, 29], algorithm evolution [30, 31, 32, 33], reward design [34], and Neural Architecture Search (NAS) [35, 36]. Notably, LLMs play a role in sampling new solutions. However, their optimization strategies depend on externally introduced natural selection mechanisms and are less effective in numerical optimization scenarios [37]. LLaMoCo [38] and EoH [39] use LLM to generate code to solve optimization problems, but the performance of LLaMoCo depends on carefully designed instructions and prompts, and EoH has expensive evaluation costs. TNPs [40], ExPT [41] and LICO [42] use transformer structures to solve the BBO problem and have achieved good results. However, TNPs requires contextual information of the target problem, and neither ExPT nor LICO can be directly used to solve tasks with different dimensions from the training task. These methods lack the universal applicability as pretrained BBO models due to a deficiency in generating capabilities across tasks.
All the above methods cannot be the zero-shot optimizer. The first two categories need to adjust the hyperParameters when optimizing the new tasks, while the latter must fine-tune the instructions to achieve satisfactory results.
3 Pretrained Optimization Model
@@ 3.1 Problem Definition
A black-box optimization problem can be transformed as a minimization problem, and constraints may exist for corresponding solutions: $\min\limits_{\mathbf{x}}\ f(\mathbf{x}),s.t.\ x_{i}\in[l_{i},u_{i}]$, where $\mathbf{x}=(x_{1},x_{2},\cdots,x_{d})$ represents the solution of optimization problem $f$, the lower and upper bounds $\mathbf{l}=(l_{1},l_{2},\cdots,l_{d})$ and $\mathbf{u}=(u_{1},u_{2},\cdots,u_{d})$, and $d$ is the dimension of $\mathbf{x}$. For more background information on evolutionary algorithms, see Appendix A.
Definition 1 Zero-shot Optimization. Zero-shot optimization refers to an optimizer that is applied directly to solve a continuous black-box optimization problem $f$ without any tuning. This means that the optimizer does not require any contextual information about $f$ and can be directly used to handle problems of any dimensionality.
Definition 2 Few-shot Optimization. Alternatively, it is permissible to fine-tune the optimizer using a small portion of the function evaluation budget for the objective task, and then use the fine-tuned optimizer to solve $f$.
3.2 Classic Population Optimization Algorithm
In this section, we use Differential Evolution (DE) as an example to review classic evolutionary algorithms. DE [20, 43] is a prominent family within evolutionary algorithms (EAs), known for its advantageous properties such as rapid convergence and robust performance [44, 45]. The optimization strategy of DE primarily involves mutation and crossover operations.
The classic DE/rand/1 crossover operator is illustrated in Eq. (1) (additional examples are listed in Appendix A.2). Each mutation strategy can be viewed as a specific instance of Eq. (2); Further details are provided in Appendix A.2. Additionally, we represent the mutation strategy in a matrix form, as shown in Eq. (3). The matrix $\mathbf{S}$ evolves with the generation index $t$, indicating that the mutation strategy adapts across different generations. Consequently, we propose a module to enhance the performance of the mutation operation, which leverages the information from the population of the $t$th generation to generate $\mathbf{S}^{t}$. This serves as the motivation for our design of the LMM.
In the crossover phase at step $t$, DE uses a fixed crossover probability $cr_{i}^{t}\in[0,1]$ for each individual $\mathbf{x}_{i}^{t}$ in the population, as shown in Eq. (9). The crossover strategy for the entire population can then be expressed as a vector $\mathbf{cr}^{t}=(cr_{1}^{t},cr_{2}^{t},\cdots,cr_{N}^{t})$. Our goal is to design a module that adaptively generates $\mathbf{cr}^{t}$ using the information from the population. This approach allows for the automatic design of the crossover strategy by controlling the parameter $cr$. This serves as the motivation for our design of LCM.
3.3 Design of POM
A population consists of $n$ individuals, denoted as $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{n}\}$. In this paper, $\mathbf{X}$ is also treated as $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{n}]^{T}$ to support matrix operations. We feed POM an initial random population $\mathbf{X}^{0}$ at step 0, specify the evolution generation $T$ for it, and hope that it can generate a population $\mathbf{X}^{T}$ close to the global optimum at step $T$, as shown in $\mathbf{X}^{T}=POM(\mathbf{X}^{0},T|\theta)$, where $\theta\in\Omega$ is the parameters of POM, where $\Omega$ stands for the strategy space. The goal of training POM is to find an optimal $\theta$ in $\Omega$. As shown in Fig. 1, POM consists of LMM, LCM and SM.
LMM
LMM generates candidate solutions $\mathbf{v}_{i}^{t}$ for individual $\mathbf{x}_{i}^{t}$ through Eq. (2), which enables the population information to be fully utilized in the process of generating candidate solutions $\mathbf{v}_{i}^{t}$.
$\mathbf{v}_{i}^{t}=\sum_{j}^{N}{w_{i,j}\mathbf{x}_{j}^{t}}\quad(\forall w_{i,j}\in\mathbb{R},w_{i,i}\neq 0)$ (2)
Further, we organize Eq. (2) into a matrix form, as shown in Eq. (3).
@@ 3.3 Design of POM
$\mathbf{X}^{t}\in\mathbb{R}^{N\times d}$ is the population in generation $t$ and $\mathbf{S}^{t}\in\mathbb{R}^{N\times N}$. $\mathbf{S}$ evolves with each change in $t$, signifying a mutation strategy that adapts across generations. Consequently, it is imperative to devise a module that leverages information from the population at generation $t$ to generate $\mathbf{S}^{t}$. Any mutation operator of differential evolution, such as the classic DE/rand/1 mutation operator, can be converted into Equation (3) in the specific case of $\mathbf{S}$ (see Appendix A.2 for details). At the same time, the crossover operation of GAs can also be generalized into the form of Equation (3) [46].
The function of LMM is designed based on Multi-head self-attention (MSA) [47], as shown as follows:
@@
$\hat{\mathbf{H}}^{t}$ $=Tanh(\mathbf{H}^{t}\times\mathbf{W}_{m1}+\mathbf{b}_{m1}),\ \ \ \mathbf{Q}^{t}=Tanh(\hat{\mathbf{H}}^{t}\times\mathbf{W}_{m2}+\mathbf{b}_{m2})$ (7)
$\mathbf{K}^{t}$ $=Tanh(\hat{\mathbf{H}}^{t}\times\mathbf{W}_{m3}+\mathbf{b}_{m3}),\ \ \ \hat{\mathbf{S}}^{t}=Tanh(\frac{\mathbf{Q}^{t}\times(\mathbf{K}^{t})^{T}}{\sqrt{(d_{m})}})$
@@
The topological structure of the population significantly influences their information exchange [48]. When all individuals engage in information exchange, the algorithm’s convergence may suffer, diversity could diminish, and susceptibility to local optima increases. To address this, we introduce a mask operation during both training and testing phases, where the probability of setting each element in $\hat{\mathbf{S}}^{t}$ to 0 is $r_{mask}$. This operation enhances POM’s ability to learn efficient and robust strategies, as validated in our experiments. Consequently, $\mathbf{S}^{t}$ is derived using Eq. (8).
@@
LCM is designed based on FFN [47], as shown in Eq. (10),
@@
where $\theta_{2}=\{\mathbf{W}_{c1},\mathbf{b}_{c1},\mathbf{W}_{c2},\mathbf{b}_{c2},\mathbf{\tau}\}$ is the parameter of LCM and $\mathbf{Z}^{t}\in\mathbb{R}^{N\times 3}$ is the population information used by LCM. Here, $\mathbf{Z}^{t}=[\mathbf{z}_{1}^{t},\mathbf{z}_{2}^{t},\cdots,\mathbf{z}_{N}^{t}]$. $\mathbf{z}_{i}^{t}$ represents the relevant information of individual $\mathbf{x}^{t}_{i}$ and $\mathbf{X}^{t}$. For example, it can include the ranking information of $\mathbf{x}^{t}_{i}$, the fitness information of $\mathbf{x}^{t}_{i}$, the Euclidean distance between $\mathbf{x}^{t}_{i}$ and $\mathbf{V}^{t}_{i}$, and the distribution information of individuals within the population (such as the fitness distribution, the distance between pairs of individuals), etc. In this paper, $\mathbf{z}_{i}^{t}$ includes the following information as a case study: 1) $\hat{f}_{i}^{t}$: the normalized fitness $f(\mathbf{x}_{i}^{t})$ of $\mathbf{x}_{i}^{t}$; 2) $\hat{r}_{i}^{t}$: the centralized ranking of $\mathbf{x}_{i}^{t}$; 3) $sim_{i}^{t}$: the cosine similarity between $x_{i}^{t}$ and $v_{i}^{t}$.
$\mathbf{h}^{t}=Tanh(\mathbf{Z}^{t}\times\mathbf{W_{c1}}+\mathbf{b_{c1}}),\ \ \mathbf{\hat{h}^{t}}=layernorm(\mathbf{h}^{t}|\mathbf{\tau}),\ \ \mathbf{cr}^{t}=Sigmoid(\mathbf{\hat{h}^{t}}\times\mathbf{W_{c2}}+\mathbf{b_{c2}})$ (11)
where the activation function Sigmoid maps inputs to the range $(0,1)$. $\mathbf{W_{c1}}\in\mathbb{R}^{3\times d_{c}}$, $\mathbf{W_{c2}}\in\mathbb{R}^{d_{c}\times 1}$, $\mathbf{\tau}$ is the learnable parameters of layernorm [49]. $\mathbf{b_{c1}}$ and $\mathbf{b_{c2}}$ are vectors with dimensions $d_{c}$ and $1$, respectively.
Although we derive $\mathbf{cr}^{t}$ from Eq. (11) as in Eq. (9), the discrete nature of the crossover operator renders it non-differentiable, impeding gradient-based training of the LCM module. To address this limitation, we introduce the gumbel_softmax method [50], providing an efficient gradient estimator that replaces non-differentiable samples from a categorical distribution with differentiable samples from a novel Gumbel-Softmax distribution.
@@
$\mathbf{r}_{i}^{t}$ $=rand(d),\ \ \ \mathbf{cv}_{i}^{t}=gumbel\_softmax(cat(\mathbf{r}_{i}^{t},tile(cr_{i}^{t},d))),$ (12)
$\mathbf{u}_{i}^{t}$ $=\mathbf{cv}_{i,0}^{t}\cdot\mathbf{x}_{i}^{t}+\mathbf{cv}_{i,1}^{t}\cdot\mathbf{v}_{i}^{t},\ \ \ \mathbf{U}^{t}=[\mathbf{u}_{1}^{t},\mathbf{u}_{2}^{t},\cdots,\mathbf{u}_{N}^{t}]$
@@
(a) POM
(b) training
(c) testing
Figure 1: In the figure, $\mathbf{X}^{0}$ is the initial random population. (a) The overall architecture of the POM. (b) POM training process. Here $T$ is the size of the inner loop iteration step during training, and the training function should be differentiable. (c) POM testing process. Here, $T$ is the number of iterations of the testing process and $f$ is the target task. $f$ does not have to be differentiable. Here we directly apply the trained POM to solve $f$ without requiring gradient information.